# Towards Heterogeneous Continual Graph Learning via Meta-knowledge Distillation

### v1

@@ C.1 Additional Ablation Study
Figure 5: Average performance (AP) and average forgetting (AF) of the MKD method under three settings: with logit-level distillation, with semantic-level distillation, and with both.
@@
C.2 Additional Hyperparameters Analysis
@@
Figure 6: Sensitivity analysis of the logit-level distillation weight $\lambda_{logit}$ in the HKD module. Left: Average Performance (AP); Right: Average Forgetting (AF).
@@
C.3 Forgetting-aware Gap
$\text{FaG}=a_{T}^{FT}-a_{T}^{CL}$ (18)
Table 4: Forgetting-aware Gap (FaG) comparison on four datasets with HAN as the backbone model (averaged over 5 runs). Lower is better.
@@ C.3 Forgetting-aware Gap
MKD (Ours) 0.63 1.31 9.90 6.88
Results Analysis: The experimental results indicate that the Joint Train method shows consistently large Final-task Performance Gaps (FaG) across all benchmark datasets, suggesting that directly replaying all data severely impairs adaptability to new tasks. MetaCLGraph and FTF-ER also suffer high FaG values on Freebase and Yelp (13.68 and 10.59), reflecting performance degradation under distribution shifts—a manifestation of the stability-plasticity dilemma. ER-GNN achieves the lowest FaG (0.22) on DBLP, showing good adaptability, but its FaG sharply rises to 8.46 on Yelp, indicating that simple replay mechanisms struggle with complex heterogeneous structures. In contrast, our MKD method consistently maintains lower FaG values. Notably, it achieves 1.31 on IMDB and 6.88 on Yelp, outperforming experience-replay and regularization-based methods like TWP (2.27 on IMDB). These results validate that MKD, through its meta-learning and heterogeneity-aware distillation design, effectively mitigates final-task degradation while preserving task adaptability. Overall, FaG serves as a useful metric to quantify the trade-off between knowledge retention and adaptation, further supporting the effectiveness of MKD in heterogeneous continual learning.
(a) Yelp (3-way)
(b) Yelp (2-way)
Figure 7: Visualization: Accuracy matrices on the Yelp dataset under both 2-way (a) and 3-way (b) settings.
Appendix D Limitations
This work is confined to domain-incremental setting and primarily addresses the node classification task. In future work, we plan to extend our approach to other settings (e.g., class-incremental) and broader graph learning tasks, such as link prediction in recommendation systems.
Appendix E Broader Impacts
Heterogeneous graphs are widely used to model complex relational systems in critical domains such as biomedical networks, social platforms, and recommender systems. Our work on Heterogeneous Continual Graph Learning (HCGL) contributes to improving the robustness and adaptability of graph models under dynamic environments. This can benefit downstream applications such as real-time recommendation, dynamic knowledge graph completion, and biomedical discovery, where information constantly evolves. By enabling efficient knowledge retention and adaptation across tasks, our framework promotes safer and more reliable deployment of graph-based models in real-world systems.

### v2

@@ C.2. Additional Ablation Study
Figure 7. Average performance (AP) and average forgetting (AF) of the HERO method under three settings: with logit-level distillation, with semantic-level distillation, and with both.
C.3. Additional Hyperparameters Analysis
@@
Figure 8. Sensitivity analysis of the logit-level distillation weight $\lambda_{logit}$ in the HKD module. Left: Average Performance (AP); Right: Average Forgetting (AF).
@@
Table 6. Forgetting-aware Gap (FaG) comparison on four datasets with HAN as the backbone model (averaged over 5 runs). Lower is better.
@@
HERO (Ours) 0.63 1.31 9.90 6.88
C.4. Forgetting-aware Gap
(18) $\text{FaG}=a_{T}^{FT}-a_{T}^{CL}$
Results Analysis The experimental results indicate that the Joint Train method shows consistently large Final-task Performance Gaps (FaG) across all benchmark datasets, suggesting that directly replaying all data severely impairs adaptability to new tasks. MetaCLGraph and FTF-ER also suffer high FaG values on Freebase and Yelp (13.68 and 10.59), reflecting performance degradation under distribution shifts—a manifestation of the stability-plasticity dilemma. ER-GNN achieves the lowest FaG (0.22) on DBLP, showing good adaptability, but its FaG sharply rises to 8.46 on Yelp, indicating that simple replay mechanisms struggle with complex heterogeneous structures. In contrast, our HERO method consistently maintains lower FaG values. Notably, it achieves 1.31 on IMDB and 6.88 on Yelp, outperforming experience-replay and regularization-based methods like TWP (2.27 on IMDB). These results validate that HERO, through its meta-learning and heterogeneity-aware distillation design, effectively mitigates final-task degradation while preserving task adaptability. Overall, FaG serves as a useful metric to quantify the trade-off between knowledge retention and adaptation, further supporting the effectiveness of HERO in heterogeneous continual learning.